### v1

@@ Abstract
To solve many problems on graphs, graph traversals are used, the usual variants of which are the depth-first search and the breadth-first search. Implementing a graph traversal we consequently reach all vertices of the graph that belong to a connected component. The breadth-first search is the usual choice when constructing efficient algorithms for finding connected components of a graph. Methods of simple iteration for solving systems of linear equations with modified graph adjacency matrices and with the properly specified right-hand side can be considered as graph traversal algorithms. These traversal algorithms, generally speaking, turn out to be non-equivalent neither to the depth-first search nor the breadth-first search. The example of such a traversal algorithm is the one associated with the Gauss-Seidel method. For an arbitrary connected graph, to visit all its vertices, the algorithm requires not more iterations than that is required for BFS. For a large number of instances of the problem, fewer iterations will be required.
@@ Introduction
Many problems from applications related to infrastructure reliability (transport networks, data networks, large integrated circuits, etc.) may be formulated as graph connectivity problems. These are problems such as finding connected components of a graph, finding its articulation points, bridges, etc.
Let $G$ be a simple labeled graph, that is, an undirected graph without multiple edges and loops, $V$ is the set of its vertices, $E$ is the set of its edges, which have cardinalities $n$ and $m$, respectively. We will assume that the vertices of the graph are labeled (numbered) in some, generally speaking, arbitrary order from $1$ to $n$. A connected component of an undirected graph is a connected subgraph that is not part of any larger connected subgraph. The connected component is defined by the set of its vertices.
Graph traversal is the iterative process during which, begining from some starting vertex, transitions along the edges of the graph are implemented. The traversal completed if all vertices of the connected component have been visited. By implementing graph traversal by an algorithm, we mean the iterative process of obtaining the vertices through which the sequential transitions along graph edges are carried out.
@@
The depth-first search (DFS) strategy [1] is to go as deep into a graph as possible. The search algorithm is described recursively: we go through all the edges incident to the current vertex; if an edge leads to a vertex that has not been visited previously, then we run the algorithm from this unconsidered vertex, and then return and continue to iterate over the edges. Return occurs if there are no edges left, which is incident to the current vertex and that lead to yet unvisited vertices of the connected component.
Usually, the breadth-first search (BFS) [1] used to find connected components computationally efficiently for large graphs. Implementing BFS, begining from a starting vertex that belongs to some connected component, we perform transitions to yet unvisited vertices that adjacent to the vertices visited at the previous iteration. At the process of such iterations, we construct reachability tree for the connected component of a graph. At the $(k\!+\!1)$-th level of this tree, there are vertices, which are reached from the starting vertex after sequential transitions along $k$ edges of some simple chain in the graph. The construction of the reachability tree for a connected component can be done in the worst case in at least $\ell_{\max}$ iterations, where $\ell_{\max}$ is the length of the simple chain, which connects starting vertex to the vertex farthest from it.
BFS was first used by K. Zuse in 1945, but was not published until 1972 [2]. Subsequently, this algorithm was rediscovered in 1959 by E. Moore as an algorithm for finding the shortest path in a labyrinth [3] and was later independently rediscovered by Lee in the context of the problem of routing conductors on printed circuit boards [4].
@@
where $x^{(0)}\!=\!e_{s}$, $e_{s}$ is the $s$-th unit vector of the standard basis in $\mathbb{R}^{n}$, i.e., all of its components are zero except the $s$-th component. At each iteration, some components of the multiplied vector become non-zero. The indices of such components of $x^{(k)}$ are numbers of the vertices visited after the $k$-th iteration of algebraic BFS.
Algebraic BFS is implemented in software libraries designed to most effectively leverage the features of modern computer architectures for parallel computing and memory optimization. They are focused on working with sparse graphs, which are most often found in applications with large graphs. Such low-level implementations in practice can be faster than implementations of theoretically optimal combinatorial BFS [5, 6, 7, 8, 9, 10]. Since computational complexity of BFS is of $O(m\!+\!n)$, for sparse graphs, that is those for which $m\!=\!O(n)$, it allows to obtain algorithms with linear computational complexity [11], which is the best possible for the problem.
And although matrix computations can be efficiently parallelized for computing of the current level of reachability tree, due to its inevitable sequential nature, there are no sublinear implementations of BFS, i.e., there is no implementations with computational complexity of $O(n^{c})$, where $10\!<\!c\!<\!$ 1.
@@
When implementing a combinatorial approach, beginning from the starting vertex, the possible options for moving along the edges of a graph are consequently considered, and then we do transitions to yet unreached vertices along some selected edges.
Combinatorial traversal algorithms, being an asymptotically optimal, in practice may be too slow for solving real applied problems with large graphs. The main part of the algorithms is a combinatorial search on possible options for transitions from the current vertex. After this search is performed, we do the transitions. But it turn out to be too slow in comparison to linear-algebaic transformation (1) that gives all possible transitions at once and can be implemented in parallel.
@@
When implementing an algebraic approach, at each iteration of the graph traversal we do some linear transformation of the state vector $x^{(k)}$ to perform an iteration of the traversal. Analysis of components of the state vector allows us to determine the vertices visited after the iteration has been completed. A visit to a vertex $i\!\in\!V$ is registered if the value of the $i$-th state vector’s component ceases to be zero after the multiplication (1). An example of algebraic approach is algebraic BFS, which uses the operation of multiplying the state vector by adjacency matrix as the linear transformation.
The graph traversal algorithms we consider use another linear transformations, different from multiplying a vector by adjacency matrix. Namely, this is a transformations implemented by a methods of simple iteration for solving a system of linear equations. Applied to the systems of linear algebraic equations (further abbreviated as SLAE) with modified graph adjacency matrices and properly specified right-hand side, variations of simple iteration methods can be considered as graph traversal algorithms. These numerical methods may provide two ways for implementing graph traversal, one of which is performed by the Jacobi method and it is equivalent to BFS, and the other performed by the Gauss-Seidel method and it gives a traversal algorithm, which is not equivalent to BFS.
Unlike DFS and BFS, applied to a graph traversal, Gauss-Seidel iterations leverage given vertex numbering in a graph. Numbering (labeling) of vertices is a mandatory parameter of a problem instance that cannot be ignored performing computations. We call a simple chain in graph as correct chain if it is a chain with ascending order of numbers of vertices in it. The approach implemented by the Gauss-Seidel method gives a traversal algorithm, which is neither DFS nor BFS. At each iteration, at first, one BFS iteration is performed, and then, if only these chains exist, we perform transitions through all of the vertices that belong to the correct chains, which are outgoing from already reached vertices.
@@ 1 Iterative numerical methods for solving SLAE
$x_{i}^{(k+1)}=\frac{1}{a_{ii}}\biggl{(}b_{i}-\sum\limits_{l=1}^{i-1}a_{ij}x_{j}^{(k+1)}-\sum\limits_{j=i+1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}.$ (4)4( 4 )
@@
$|a_{ii}|\geq\sum\limits_{i\neq j}|a_{ij}|$
for all $i\!=\!\overline{1,n}$, and at least one of these inequalities is strict, then iterations (3) and (4) converge to the exact solution that exists and that is unique for any right-hand side
Consider SLAE (2), where $A$ is a modified graph adjacency matrix with diagonal dominance and $b\!=\!e_{s}$. The algorithms for graph traversals that we consider further, are the ones that we obtain implementing the iterations (3) and (4). After performing not more than $n$ iterations (3) or (4) for $x^{(0)}\!=\!d\cdot e_{s}$, maybe without achieving convergence to the exact solution of the SLAE (2), we obtain a sequence of approximate solutions $x^{(k)}$. Further, rather than apprpximate solutions of a SLAE, we shall see them as a current value of the state vector $x^{(k)}$ for iteration $k$.
Transitions between vertices of a graph can be registered using these computations. At the $(k\!+\!1)$-th iteration, the transition occurs from one of the vertices reached at the $k$-th iteration to some vertex $i\!\in\!V$ if the following condition is met:
@@
That is, the vertex $i$ was not visited prior to the $(k\!+\!1)$-th iteration, and it is reached at this iteration. The frontier $\mathcal{F}^{(k+1)}$ for the $(k\!+\!1)$-th iteration is the set of the vertices that ceases to be zero at this iteration. $\mathcal{F}^{(0)}\!=\!\{s\}$.
Performing Jacobi iterations or Gauss-Seidel iterations with the given initial value of the state vector and the right-hand side of the SLAE (2), we sequentially obtain frontiers of traversals and, finally, visit all vertices of a graph.
All of the simple iteration methods for solving SLAE give the same two traversal algorithms. One of them is given by (3), and, as we shall see, it is equivalent to BFS, and the other is given by (4), and it is not equivalent to BFS.
@@ 2 Using multiplication instead of division
Moreover, it makes practical sense to use multiplying instead of division in Jacobi and Gauss-Seidel iterations to perform graph traversal. Implementations of division with floating point machine numbers are several times slower than implementations of multiplication, so, after such a replacement, the time required for implementing graph traversal using the modified iterations will be shorter.
Thus, further, considering adjacency graph matrices with initially zero diagonal elements replaced by some $d\!>\!0$, we consider iterations for the two methods as, respectively,
$x_{i}^{(k+1)}=\biggl{(}-b_{i}+\sum\limits_{j\neq i}a_{ij}x_{j}^{(k)}\biggr{)}(-d)$ (6)6( 6 )
and
$x_{i}^{(k+1)}=\biggl{(}-b_{i}+\sum\limits_{l=1}^{i-1}a_{ij}x_{j}^{(k+1)}+\sum\limits_{j=i+1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}(-d).$ (7)7( 7 )
3 BFS and CCS as combinatorial algorithms
A chain in a graph is a finite sequence of vertices in which each vertex except the last one is connected to the next one in the sequence by an edge. The length of the chain $c$ we denote as $\ell(c)$, which is the number of edges in $c$. We call a chain as a correct chain if its vertices’ labels, which are unique numbers, are given in ascending order, i.e., it is a chain $c\!=\{i_{0},i_{1},\ldots,i_{k}\}$, $i_{j}\!\in\!V$, such that $(i_{j-1},i_{j})\!\in\!E$, $i_{j-1}\!<\!i_{j}$, $j\!=\!\overline{1,k}$. Correct chains are simple chains, i.e., chains without self-intersections. A walk is a chain where edges and vertices both can be repeated in it.
@@ 3 BFS and CCS as combinatorial algorithms
The similarity and the difference between BFS and the traversal algorithm that Gauss-Seidel iterations give are as follows. As for BFS, performing iterations (7), at first, we perform transitions along the edges that is incident to vertices reached at the previous iteration. The difference between these graph traversal algorithms is that if there are outgoing correct chains from the vertices visited at the current iteration, then at the same iteration, transitions through all vertices of these correct chains will be done.
This happens since, when performing the algebraic CCS, computations of individual components of the current state vector are performed using the values of the components, which have been computed earlier in the course of the current iteration. So, if vertex $j$ is adjacent to vertex $i$ and $j\!<\!i$, and the $j$-th component of the state vector ceases to be zero at the current iteration, then at the same iteration its value will be used for computing of the $i$-th component and it ceases to be a zero too if it was zero before. On the contrary, the BFS state vector is updated based only on the results of individual computations for each component of the state vector. It is good to perform these computations in parallel but it not allows the frontier to be wider than neibourghood of the previous frontier vertices. For CCS, if there are correct chains that outgo from the neibourghood of the previous frontier, we traverse through them at the same iteration.
For $(k\!+\!1)$-th iteration of the combinatorial BFS, $\mathcal{F}^{(k+1)}$ is a set of vertices reached by traversing edges, which are incident to vertices in $\mathcal{F}^{(k)}$ and which not traversed in the course of the previous $k$ iterations. Let $\mathcal{N}$ denote the set of vertices that adjacent to vertices in $\mathcal{F}^{(k)}$, $\mathcal{C}^{(k)}$ is the set of vertices in $V$ visited after $k$ iterations of traversal algorithm.
@@
Implementing iteration of CCS, in addition to traversing edges that incident to vertices in $\mathcal{F}^{(k)}$, we perform traversals of the correct chains that outgo from the vertices that reached through these incident edges. Let $C(\tilde{s},i)$ be the set of all correct simple chains outgoing from vertex $\tilde{s}\!\in\!\mathcal{N}$ and ending at vertex $i$.
@@
At step 3 of both algorithms, we implement transitions along the edges incident to the vertices from the frontier $\mathcal{F}^{(k)}$ obtained at the previous iteration. Additional transitions through correct chains from $C(\tilde{s},i)$ implemented at step 4 of the Combinatorial CCS.

### v2

@@ Abstract
To solve many problems on graphs, graph traversals are used, the usual variants of which are the depth-first search and the breadth-first search. Implementing a graph traversal we consequently reach all vertices of the graph that belong to a connected component. The breadth-first search is the usual choice when constructing efficient algorithms for finding connected components of a graph. Methods of simple iteration for solving systems of linear equations with modified graph adjacency matrices can be considered as graph traversal algorithms if we use properly specified right-hand side. These traversal algorithms, generally speaking, turn out to be non-equivalent neither to the depth-first search nor the breadth-first search. The example of such a traversal algorithm is the one associated with the Gauss-Seidel method. For an arbitrary connected graph, to visit all its vertices, the algorithm requires not more iterations than that is required for breadth-first search. For a large number of instances of the problem, fewer iterations will be required.
@@ Introduction
Many problems from applications related to infrastructure reliability (transport networks, data networks, large integrated circuits, etc.) formulated as graph connectivity problems. These are problems such as finding connected components of a graph, finding its articulation points, bridges, etc.
Let $G$ be a simple labeled graph, that is, an undirected unwheighted graph without multiple edges and loops, $V$ is the set of its vertices, $|V|\!=\!n$, $E$ is the set of its edges, $|E|\!=\!m$. We will assume that the vertices of the graph are labeled (numbered) in some, generally speaking, arbitrary order from $1$ to $n$. A connected component of an undirected graph is a connected subgraph that is not part of any larger connected subgraph. The connected component is defined by the set of its vertices.
Graph traversal is the iterative process during which, begining from some starting vertex, transitions along the edges of the graph are implemented. The traversal completed if all vertices of the connected component have been visited. By implementing graph traversal by an algorithm we mean the iterative process of obtaining the vertices through which the sequential transitions along graph edges are carried out.
@@
The depth-first search (DFS) strategy [1] is to go as deep into a graph as possible. The search algorithm is described recursively: we go through all the edges incident to the current vertex; if an edge leads to a vertex that has not been visited previously, then we run the algorithm from this unconsidered vertex, and then return and continue to iterate over the edges. Return occurs if there are no edges left which is incident to the current vertex and that lead to yet unvisited vertices of the connected component.
Usually, the breadth-first search (BFS) [1] used to find connected components computationally efficiently for large graphs. Implementing BFS, begining from a starting vertex that belongs to some connected component, we perform transitions to yet unvisited vertices that adjacent to the vertices visited at the previous iteration. At the process of such iterations, we construct reachability tree for the connected component of a graph. At the $(k+1)$-th level of this tree, there are vertices which are reached from the starting vertex after sequential transitions along $k$ edges of some simple chain in the graph. The construction of the reachability tree for a connected component can be done in $\ell_{\max}$ iterations, where $\ell_{\max}$ is the length of the shortest simple chain which connects starting vertex to the vertex farthest from it.
BFS was first used by K. Zuse in 1945, but was not published until 1972 [2]. Subsequently, this algorithm was rediscovered in 1959 by E. Moore as an algorithm for finding the shortest path in a labyrinth [3] and was later independently rediscovered by C. Lee in the context of the problem of routing conductors on printed circuit boards [4].
@@
where $x^{(0)}\!=\!e_{s}$, $e_{s}$ is the $s$-th unit vector of the standard basis in $\mathbb{R}^{n}$, i.e., all of its components are zero except the $s$-th component. At each iteration, some components of the multiplied vector become non-zero. The indices of such components of $x^{(k)}$ are numbers of the vertices visited after the $k$-th iteration of BFS has been implemented.
Algebraic BFS is implemented in software libraries designed to most effectively leverage the features of modern computer architectures for parallel computing and memory optimization. They are focused on working with large sparse graphs which are most often found in applications. Such low-level implementations in practice can be faster than implementations of theoretically optimal combinatorial BFS [5, 6, 7, 8, 9, 10]. Since computational complexity of BFS is of $O(m+n)$, for sparse graphs, that is those for which $m\!=\!O(n)$, it allows to obtain algorithms with linear computational complexity [11] which is the best possible for the problem.
And although computations (1) can be efficiently parallelized for computing of the current level of reachability tree, due to its inevitable sequential nature, there are no sublinear implementations of BFS, i.e., there is no implementations with computational complexity of $O(n^{c})$, where $10\!<\!c\!<\!$ 1.
@@
When implementing a combinatorial approach, beginning from the starting vertex, the possible options for moving along the edges of a graph are considered consequently, and then we do transitions to yet unreached vertices along some selected edges.
Combinatorial traversal algorithms, being asymptotically optimal, in practice may be too slow for solving real applied problems with large graphs. The main part of the algorithms is a combinatorial search on possible options for transitions from the current vertex. After this search is performed, we do the transitions. But it turn out to be too slow in comparison to linear-algebaic transformation (1) that gives all possible transitions at once and can be implemented in parallel.
@@
When implementing an algebraic approach, at each iteration of the graph traversal we do some linear transformation of the state vector $x^{(k)}$ to perform an iteration of the traversal. Analysis of components of the state vector allows us to determine the vertices visited after the iteration has been completed. A visit to a vertex $i\!\in\!V$ is registered if the value of the $i$-th state vector’s component ceases to be zero after this linear transformation. An example of algebraic approach is algebraic BFS which uses the operation of multiplying the state vector by adjacency matrix (1) as the linear transformation.
The graph traversal algorithms we consider use another linear transformations, different from multiplying a vector by adjacency matrix. Namely, this is a transformations implemented by a methods of simple iteration for solving a system of linear equations. Applied to the systems of linear algebraic equations (further abbreviated as SLAE) with modified graph adjacency matrices and properly specified right-hand side, variations of simple iteration methods can be considered as graph traversal algorithms. These numerical methods may provide two ways for implementing graph traversal, one of which is performed by the Jacobi method and it is equivalent to BFS, and the other performed by the Gauss-Seidel method and it gives a traversal algorithm which is not equivalent to BFS.
Unlike DFS and BFS, applied to a graph traversal, Gauss-Seidel iterations leverage given vertex numbering in a graph. Numbering (labeling) of vertices is a mandatory parameter of a problem instance that cannot be ignored performing computations. We call a simple chain in graph as correct chain if it is a chain with ascending order of numbers of vertices in it. Leveraging these chains, the approach implemented by the Gauss-Seidel method gives a traversal algorithm which is neither DFS nor BFS. At each iteration, at first, one BFS iteration is performed, and then, if only these chains exist, transitions made along edges of the correct chains which are outgoing from the new reached at this iteration vertices.
@@ 1 Iterative numerical methods for solving SLAE
$x_{i}^{(k+1)}=\frac{1}{a_{ii}}\biggl{(}b_{i}-\sum\limits_{j=1}^{i-1}a_{ij}x_{j}^{(k+1)}-\sum\limits_{j=i+1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}.$ (4)4( 4 )
@@
$|a_{ii}|=d\geq\sum\limits_{i\neq j}|a_{ij}|$
for all $i\!=\!\overline{1,n}$, and at least one of these inequalities is strict, then approximate solutions obtained at iterations (3) and (4) converge to the exact solution that exists and that is unique for any right-hand side $b$.
Consider SLAE (2), where $A$ is a modified graph adjacency matrix with diagonal dominance and $b\!=\!e_{s}$. The graph traversals algorithms that we consider further, are the ones that we obtain implementing the iterations (3) and (4). After performing not more than $n$ iterations (3) or (4) with $x^{(0)}\!=\!d\cdot e_{s}$ as initial vector, maybe without achieving convergence to the exact solution of the SLAE (2), we obtain a sequence of approximate solutions $x^{(k)}$. Further, rather than approximate solutions of a SLAE, we shall consider $x^{(k)}$ as current value of the state vector $x^{(k)}$ for iteration $k$.
Transitions between vertices of a graph can be registered using these computations. At the $(k+1)$-th iteration, the transition occurs to some vertex $i$ from one of the vertices reached at the $k$-th iteration if the following condition is met:
@@
That is, the vertex $i$ was not visited prior to the $(k+1)$-th iteration, and it is reached at this iteration. The frontier $\mathcal{F}^{(k+1)}$ for the $(k+1)$-th iteration is the set of the vertices for which corresponding components of the state vector cease to be zero at this iteration. $\mathcal{F}^{(0)}\!=\!\{s\}$.
Performing Jacobi iterations or Gauss-Seidel iterations with the given initial value of the state vector and the right-hand side of the SLAE (2), we sequentially obtain frontiers of traversals and, finally, visit all vertices of a connected graph.
All of the simple iteration methods for solving SLAE give the same two traversal algorithms in the sence of the sequentially obtained traversals. One of them is given by (3), and, as we shall see, it is equivalent to BFS, and the other is given by (4), and it is not equivalent to BFS.
@@ 2 Using multiplication instead of division
Moreover, it makes practical sense to use multiplying instead of division in Jacobi and Gauss-Seidel iterations to perform graph traversal. Implementations of division with floating point machine numbers are several times slower than implementations of multiplication. So, after such a replacement, the time required for implementing graph traversal using the modified iterations will be shorter.
Thus, further, considering adjacency graph matrices with initially zero diagonal elements replaced by some $d\!>\!0$, we consider iterations for the two methods as, respectively, (6) and (7):
$x_{i}^{(k+1)}=\biggl{(}-b_{i}+\sum\limits_{j\neq i}a_{ij}x_{j}^{(k)}\biggr{)}(-d),$ (6)6( 6 )
$x_{i}^{(k+1)}=\biggl{(}-b_{i}+\sum\limits_{j=1}^{i-1}a_{ij}x_{j}^{(k+1)}+\sum\limits_{j=i+1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}(-d).$ (7)7( 7 )
The value of $d$ can be arbitrary positive value.
3 Combinatorial graph traversal algorithms
A chain in a graph is a finite sequence of vertices in which each vertex except the last one is connected to the next one in the sequence by an edge. A chain may be considered as sequence of these edges. The length of the chain $c$ we denote as $\ell(c)$ which is the number of edges in $c$. A chain with the same initial and last vertex is called a cycle. We call a chain as a correct chain if its vertices’ labels which are unique numbers, are given in ascending order, i.e., it is a chain $c\!=\{i_{0},i_{1},\ldots,i_{k}\}$, $i_{j}\!\in\!V$, such that $(i_{j-1},i_{j})\!\in\!E$, $i_{j-1}\!<\!i_{j}$, $j\!=\!\overline{1,k}$. Correct chains are simple chains, i.e., chains without self-intersections. A walk is a chain where vertices and edges both can be repeated in it.
@@ 3 Combinatorial graph traversal algorithms
The similarity and the difference between BFS and CCS are as follows. As for BFS, performing iterations (7), at first, we perform transitions along the edges that is incident to vertices reached at the previous iteration. The difference between these graph traversal algorithms is that if there are outgoing correct chains from the vertices visited at the current iteration, then at the same iteration, transitions through all vertices of these correct chains will be done.
This happens since, when performing the iterations (7), computations of individual components of the current state vector are performed using the values of the components which have been computed earlier in the course of the current iteration. So, if vertex $j$ is adjacent to vertex $i$ and $j\!<\!i$, and the $j$-th component of the state vector ceases to be zero at the current iteration, then at the same iteration its value will be used for computing of the $i$-th component and it ceases to be a zero too if it was zero before. On the contrary, the BFS state vector is updated based only on the results of individual computations for each component of the state vector. It is good since it allows to perform these computations in parallel, but it not allows the frontier to be wider than neibourghood of the previous frontier vertices. For CCS, if there are correct chains that outgo from the neibourghood of the previous frontier, we traverse through them at the same iteration adding new vertices to the frontier.
For $(k+1)$-th iteration of the combinatorial BFS, $\mathcal{F}^{(k+1)}$ is a set of vertices reached by traversing edges which are incident to vertices in $\mathcal{F}^{(k)}$ and which not traversed in the course of the previous $k$ iterations. Let $\mathcal{N}$ be this set. $\mathcal{C}^{(k)}$ is the set of vertices in $V$ visited after $k$ iterations of traversal algorithm.
@@
Implementing iteration of CCS, in addition to traversing edges that incident to vertices in $\mathcal{F}^{(k)}$, we perform traversals of the correct chains that outgo from the vertices that reached through these incident edges. Let $C(\tilde{s},i)$ be the set of all correct chains outgoing from vertex $\tilde{s}\!\in\!\mathcal{N}$ and ending at vertex $i$.
@@
So, at the step 3 of both algorithms, we implement transitions along the edges incident to the vertices from the frontier $\mathcal{F}^{(k)}$ obtained at the previous iteration. Additional transitions through correct chains from $C(\tilde{s},i)$ for $\tilde{s}\!\in\!\mathcal{N}$ implemented at the step 4 of the Combinatorial CCS.